# Mean-field limits for Consensus-Based Optimization and Sampling

### v2

@@ 2.4 Quantitative mean-field limit for CBO
$\mathbf{E}\left[\sup_{s\in[0,t]}\bigl{\lvert}X_{s}^{(j)}-\overline{X}_{s}^{(j)}\bigr{\rvert}^{p}\right]$ $\leqslant 2^{p-1}\mathbf{E}\left(\int_{0}^{t}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert\,\mathrm{d}s\right)^{p}$
$\qquad+C_{\rm BDG}2^{p-1}\mathbf{E}\left(\int_{0}^{t}\left\lVert\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert_{\rm F}^{2}\,\mathrm{d}s\right)^{\frac{p}{2}}$
$\leqslant(2T)^{p-1}\mathbf{E}\int_{0}^{t}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s$
$\qquad+\,C_{\rm BDG}2^{p-1}T^{\frac{p}{2}-1}\mathbf{E}\int_{0}^{t}\left\lVert\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert_{\rm F}^{p}\,\mathrm{d}s.$ (2.13)
@@
$\mathbf{E}\int_{0}^{t}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s$ $\leqslant 2^{p-1}\mathbf{E}\int_{0}^{t}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)\,\right\rvert^{p}\mathrm{d}s$
$\qquad+2^{p-1}\mathbf{E}\int_{0}^{t}\left\lvert b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s.$ (2.14)
@@
$\mathbf{E}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)\right\rvert^{p}$ $=\mathbf{E}\left\lvert\Bigl{(}\overline{X}^{(j)}_{s}-X^{(j)}_{s}+\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\Bigr{)}\right\rvert^{p}$
$\leqslant 2^{p-1}\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+2^{p-1}\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}.$ (2.15)
@@
$\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}$ $=\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega\setminus\Omega_{J,s}}\right]$
$\qquad+\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega_{J,s}}\right].$ (2.17)
@@
    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega\setminus\Omega_{J,s}}\right]\leqslant C\mathbf{E}\left[W_{p}\left(\mu^{J}_{s},\overline{\mu}^{J}_{s}\right)^{p}\right]$
@@
    $\mathbf{E}\left[W_{p}\left(\mu^{J}_{s},\overline{\mu}^{J}_{s}\right)^{p}\right]\leqslant\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\left\lvert X_{s}^{(j)}-\overline{X}^{(j)}_{s}\right\rvert^{p}\right]=\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p},$
@@
    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega\setminus\Omega_{J,s}}\right]\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}.$
  • We now bound the second term in (2.17). From Item (b) (see Lemma 3.5) and (2.4) we have
    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{q}\right]\leqslant 2^{q-1}\left(\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)\right\rvert^{q}+\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{q}\right)\leqslant 2^{q}\kappa.$
    Therefore, we obtain by Hölder’s inequality and Lemma 2.5 that
    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega_{J,s}}\right]\leqslant\left(\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{q}\right]\right)^{\frac{p}{q}}\left(\mathbf{P}\left[\Omega_{J,s}\right]\right)^{{}^{\frac{q-p}{q}}}\leqslant 2^{p}\kappa^{\frac{p}{q}}J^{-\frac{q-p}{2p}}.$
@@
$\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+CJ^{-\frac{q-p}{2p}},$
@@
$\mathbf{E}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+CJ^{-\frac{q-p}{2p}}.$
@@
$\mathbf{E}\left\lvert b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}=\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\rho}_{s}\right)\right\rvert^{p}\leqslant CJ^{-\frac{p}{2}}.$
@@ Conclusion
$\mathbf{E}\left[\int_{0}^{t}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s\right]\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{q-p}{2p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
@@
$\mathbf{E}\left[\int_{0}^{t}\left\lVert\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert^{p}\,\mathrm{d}s\right]\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{q-p}{2p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
@@
The conclusion then follows from Grönwall’s lemma, since $\frac{q-p}{2p^{2}}\leqslant\frac{q-(2\vee p_{\mathcal{M}})}{2(2\vee p_{\mathcal{M}})^{2}}$ for the value of $p$ considered. ∎
@@
Remark 2.10.
A crucial detail in the proof is that Item (a) only assumes that $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p}$ and not $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p,R}$; see Corollary 3.3 below. This allows us to define $\Omega_{J,t}$ in (2.16) only in terms of the i.i.d. mean-field particles $\overline{X}^{(j)}_{t}$ for which Lemma 2.5 can be applied; the event $\frac{1}{J}\sum_{j=1}^{J}\left\lvert X^{(j)}_{t}\right\rvert^{p}\geqslant R$ need not be included in the definition of $\Omega_{J,t}$
Remark 2.11.
In his famous lecture notes [MR1108185], Sznitman shows for a laboratory example that
$\forall j\in\llbracket 1,J\rrbracket,\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\leqslant CJ^{-\frac{p}{2}}.$
More precisely, this statement is proved for $p=1$, and the author notes that the result generalizes to higher moment, which can be achieved by using, among other technical adaptations of his argument, the Marcinkiewicz–Zygmund inequality [MR1476912]. Although CBO does not fit within Sznitman’s simplified framework, Item (c) enables to show a bound with the same rate for the second term on the right-hand side of (2.14),
$\mathbf{E}\biggl{[}\left\lvert b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega\setminus\Omega_{J,t}}\biggr{]}\leqslant\frac{C}{J^{\frac{p}{2}}}.$
The limiting factor, for the overall rate of convergence in Theorem 2.6, is the bound on the probability $\mathbf{P}[\Omega_{J,t}]$.

### v3

@@ 2.4 Quantitative mean-field limit for CBO
$\mathbf{E}\left[\sup_{s\in[0,t]}\bigl{\lvert}X_{s}^{(j)}-\overline{X}_{s}^{(j)}\bigr{\rvert}^{p}\right]$ $\leqslant 2^{p-1}\mathbf{E}\left(\int_{0}^{t}\left\lvert b\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\rho}_{s}\Bigr{)}\right\rvert\,\mathrm{d}s\right)^{p}$
$\qquad+C_{\rm BDG}2^{p-1}\mathbf{E}\left(\int_{0}^{t}\left\lVert\sigma\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-\sigma\Bigl{(}\overline{X}_{s}^{(j)},\overline{\rho}_{s}\Bigr{)}\right\rVert_{\rm F}^{2}\,\mathrm{d}s\right)^{\frac{p}{2}}$
$\leqslant(2T)^{p-1}\mathbf{E}\int_{0}^{t}\left\lvert b\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\rho}_{s}\Bigr{)}\right\rvert^{p}\,\mathrm{d}s$
$\qquad+\,C_{\rm BDG}2^{p-1}T^{\frac{p}{2}-1}\mathbf{E}\int_{0}^{t}\left\lVert\sigma\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-\sigma\Bigl{(}\overline{X}_{s}^{(j)},\overline{\rho}_{s}\Bigr{)}\right\rVert_{\rm F}^{p}\,\mathrm{d}s.$ (2.13)
@@
$\mathbf{E}\int_{0}^{t}\left\lvert b\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\rho}_{s}\Bigr{)}\right\rvert^{p}\,\mathrm{d}s$ $\leqslant 2^{p-1}\mathbf{E}\int_{0}^{t}\left\lvert b\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\Bigr{)}\,\right\rvert^{p}\mathrm{d}s$
$\qquad+2^{p-1}\mathbf{E}\int_{0}^{t}\left\lvert b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\rho}_{s}\Bigr{)}\right\rvert^{p}\,\mathrm{d}s.$ (2.14)
@@
$\mathbf{E}\left\lvert b\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\Bigr{)}\right\rvert^{p}$ $=\mathbf{E}\left\lvert\Bigl{(}\overline{X}^{(j)}_{s}-X^{(j)}_{s}+\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\Bigr{)}\right\rvert^{p}$
$\leqslant 2^{p-1}\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+2^{p-1}\mathbf{E}\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}.$ (2.15)
@@
$\mathbf{E}\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}$ $=\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}\mathsf{1}_{\Omega\setminus\Omega_{J,s}}\right]$
$\qquad+\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}\mathsf{1}_{\Omega_{J,s}}\right].$ (2.17)
@@
    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}\mathsf{1}_{\Omega\setminus\Omega_{J,s}}\right]\leqslant C\mathbf{E}\left[W_{p}\Bigl{(}\mu^{J}_{s},\overline{\mu}^{J}_{s}\Bigr{)}^{p}\right]$
@@
    $\mathbf{E}\left[W_{p}\Bigl{(}\mu^{J}_{s},\overline{\mu}^{J}_{s}\Bigr{)}^{p}\right]\leqslant\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\left\lvert X_{s}^{(j)}-\overline{X}^{(j)}_{s}\right\rvert^{p}\right]=\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p},$
@@
    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}\mathsf{1}_{\Omega\setminus\Omega_{J,s}}\right]\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}.$
  • We now bound the second term in (2.17). From Item (b) (see Lemma 3.5) as well as (2.4) and Proposition A.3 we have
    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{q}\right]\leqslant 2^{q-1}\left(\mathbf{E}\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}\right\rvert^{q}+\mathbf{E}\left\lvert\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{q}\right)\leqslant 2^{q-1}\left(\kappa+\widetilde{\kappa}\right)$
    where $\widetilde{\kappa}:=C_{2}\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{q}\right]<\infty$. Here, $C_{2}>0$ is the constant from Proposition A.3. Therefore, we obtain by Hölder’s inequality and Lemma 2.5 that
    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}\mathsf{1}_{\Omega_{J,s}}\right]\leqslant\left(\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{q}\right]\right)^{\frac{p}{q}}\left(\mathbf{P}\left[\Omega_{J,s}\right]\right)^{{}^{\frac{q-p}{q}}}\leqslant 2^{\frac{(q-1)p}{q}}\left(\kappa+\widetilde{\kappa}\right)^{\frac{p}{q}}J^{-\frac{q-p}{2p}}.$
@@
$\mathbf{E}\left\lvert\mathcal{M}_{\beta}\bigl{(}\mu^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}\right\rvert^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+CJ^{-\frac{q-p}{2p}},$
@@
$\mathbf{E}\left\lvert b\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\Bigr{)}\right\rvert^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+CJ^{-\frac{q-p}{2p}}.$
@@
$\mathbf{E}\left\lvert b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\rho}_{s}\Bigr{)}\right\rvert^{p}=\mathbf{E}\left\lvert\mathcal{M}_{\beta}\bigl{(}\overline{\mu}^{J}_{s}\bigr{)}-\mathcal{M}_{\beta}\bigl{(}\overline{\rho}_{s}\bigr{)}\right\rvert^{p}\leqslant CJ^{-\frac{p}{2}}.$
@@ Conclusion
$\mathbf{E}\left[\int_{0}^{t}\left\lvert b\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-b\Bigl{(}\overline{X}_{s}^{(j)},\overline{\rho}_{s}\Bigr{)}\right\rvert^{p}\,\mathrm{d}s\right]\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{q-p}{2p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
@@
$\mathbf{E}\left[\int_{0}^{t}\left\lVert\sigma\Bigl{(}X_{s}^{(j)},\mu^{J}_{s}\Bigr{)}-\sigma\Bigl{(}\overline{X}_{s}^{(j)},\overline{\rho}_{s}\Bigr{)}\right\rVert^{p}\,\mathrm{d}s\right]\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{q-p}{2p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
@@
The conclusion follows from Grönwall’s lemma, since $\frac{q-p}{2p^{2}}\leqslant\frac{q-(2\vee p_{\mathcal{M}})}{2(2\vee p_{\mathcal{M}})^{2}}$ for the value of $p$ considered. ∎
@@
A crucial detail here is the following: Item (a) only assumes that $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p}$ and not $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p,R}$; see Corollary 3.3 below. This allows us to define $\Omega_{J,t}$ in (2.16) only in terms of the i.i.d. mean-field particles $\overline{X}^{(j)}_{t}$ for which Lemma 2.5 can be applied; the event $\frac{1}{J}\sum_{j=1}^{J}\left\lvert X^{(j)}_{t}\right\rvert^{p}\geqslant R$ does not need to be included in the definition of $\Omega_{J,t}$. This trick allows to show the strong mean-field limit result (2.12), improving on the previous results from [CBO-mfl-Huang2021, CBO-hypersurface-Fornasier_2020, fornasier2021consensusbased].